OpenStudy (anonymous):

how do you find a partial fraction decomposition of -22x/ 3x^2 +7x-6

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Could you be more specific? Is it 22x/ (3x^2 +7x -6)? or is it the way you wrote it? Because this will have a huge effect on the answer :)

myininaya (myininaya):

\[\frac{-22x}{3x^2+9x-2x-6)}=\frac{-22x}{3x(x+3)-2(x+3)}=\frac{-22x}{(3x-2)(x-3)}=\frac{A}{3x-2}+\frac{B}{x-3}\]

myininaya (myininaya):

\[-22x=A(x-3)+B(3x-2)\]

myininaya (myininaya):

-22x=Ax+3Bx-3A-2B

OpenStudy (annon):

myininaya u beet me to the punch lol

myininaya (myininaya):

22x=(A+3B)x+(-3A-2B) A+3B=22 -3A-2B=0

myininaya (myininaya):

A=22-3B -3(22-3B)-2B=-66+9B-2B=-66+7B=0 7B=66 B=66/7 so A=22-3(66/7)=-44/7 so \[\frac{-22x}{3x^2+7x-6}=\frac{\frac{-44}{7}}{3x-2}+\frac{\frac{66}{7}}{x-3}=\frac{-44}{7(3x-2)}+\frac{66}{7(x-3)}\]

myininaya (myininaya):

im sorry annon i just realized u said something you can check my work if you want